hi im in 8th grade and i need help with linear equations can you help me understand them and can you tell me how to do them heres one 6(8+3a)=-2(a-4) thankyou
Answer by stanbon(75887)

6(8+3a)=-2(a-4)
48+18a=-2a+8
40 =-20a
a = -2
Cheers,
stan H.